OpenStudy (anonymous):

Peter played his favorite video game for 10 hours last week. Today Peter's parents restricted him to 5 hours a week for the next 8 weeks. Create a function, f(x), that models Peters total video game time and explain what each number in the situation represents, using complete sentences.

OpenStudy (lucaz):

i guess it could be f(x)=8x+10 x as the hours from 0 to 5 = the domain is [0,5]

OpenStudy (lucaz):

I'm not sure

OpenStudy (anonymous):

It looks right. Is there any way to be sure?

OpenStudy (lucaz):

well, I don't know =/ but at least the function makes sense

OpenStudy (anonymous):

haha well yeah. Thanks for your help, you mind helping with another question?

OpenStudy (lucaz):

yea

OpenStudy (anonymous):

Marie is logging the number of radioactive isotopes she observes every minute. She has determined to function to be f(x) = -4x + 120. Based on the situation above, describe any necessary restrictions to the domain and range. Thanks :)

OpenStudy (lucaz):

it seems f(x) must be positive, is that right?

OpenStudy (anonymous):

Yeah

OpenStudy (lucaz):

positive or greater than/equal to 0 -4x+120>=0 -4x>=-120 4x<=120 x<=30 domain would be (-∞,30] range [0,∞)

OpenStudy (anonymous):

Great thanks man :)

OpenStudy (lucaz):

sorry, but I don't know again if it's right =/

OpenStudy (anonymous):

Well, it looks ok to me

OpenStudy (lucaz):

hope so =P